OpenStudy (anonymous):

what should you substitute for y in the second equation (bottom equation) in order to solve the system by the substitution method? 2y-x=6 4x-y=8

OpenStudy (goformit100):

@jackellyn

OpenStudy (jackellyn):

2y-x = 6 2y = 6+x y = (6+x)/2 4x - ((6+x)/2) =8